For any positive number a,b,m,n satisfying the condition
m/n = a /(a+b) the difference a/b - m/n is equal to
a) nb / ma
b) mb / na
c) ma / nb
d) ab / mn

Answer by THANApHD(104)

m/n = a /(a+b)
so
n/m = (a+b)/a
na = m(a+b)
na = ma + mb - ----- -- -- -(1)

So
a/b - m/n
= (na - mb)/(nb) .. but na = ma + mb ---- (1)
= (ma + mb - mb)/(nb)
= ma/nb

ans 3